# RedTeamCUA: Realistic Adversarial Testing of Computer-Use Agents in Hybrid Web-OS Environments

### v4

@@ 8 Conclusion
9 Limitations
(3) Additionally, we observe that CUAs exhibit varying levels of vulnerability depending on the specific filename used (Appendix H.3). Although this observation alone does not justify specifying filenames in the injection, it highlights an additional consideration: filenames themselves might meaningfully influence attack outcomes, further reinforcing the need for controlled specification in our evaluations.
10 Contribution Statement
Jaylen Jones led the initial grant writing and preliminary conception of ideas for project formulation of RedTeamCUA. Jaylen Jones and Zeyi Liao led preliminary experiments to explore construction of RedTeamCUA and initial attack ideas. All primary authors contributed to adversarial attack formulation in RTC-Bench, with primary efforts coming from Zeyi Liao (e.g. contributions to high-level threat model and adversarial injection design) and Linxi Jiang (e.g. adversarial goal construction based on CIA triad). Zeyi Liao provided the benign task formulation used within RTC-Bench, significantly influencing sandbox and benchmark design. Zeyi Liao and Linxi Jiang led the main code implementation for the RedTeamCUA framework and sandbox construction. Zeyi Liao and Linxi Jiang led the implementation of and performed main experiments and ablations using the RTC-Bench benchmark. All primary authors discussed and analyzed the results. Jaylen Jones and Zeyi Liao led the primary writing of the final manuscript and appendices while all other authors provided additional suggestions and contributions. Yuting Ning helped explore and run experiments of different defense approaches and refined the writing. Eric Fosler-Lussier, Yu Su, and Zhiqiang Lin are senior authors who provide guidance and constructive suggestions throughout the project. Huan Sun is the senior lead author who conceived and oversaw the research project, contributed to the core ideas, and revised the manuscript.
11 Acknowledgment
The authors would thank colleagues from the OSU NLP group for their constructive feedback. This research was sponsored in part by NSF CAREER #1942980, NSF CNS #2112471, the Schmidt Sciences Safety Science award, and Ohio Supercomputer Center (Center, 1987). The views and conclusions contained herein are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notice herein.
@@ 11 Acknowledgment
In § 3, we describe how we construct our hybrid sandbox RedTeamCUA, by leveraging OSWorld as a backbone and integrating Docker-based web replicas from WebArena and TheAgentCompany. In§ 4, we detail the formulation of benign and adversarial goals, along with concrete examples in Appendix A. In § 5.1 and Appendix D, we provide details on the evaluated CUAs, evaluation metrics (i.e., SR, ASR and AR) and AWS configuration. Upon acceptance, we will open-source all our related materials, including our sandbox RedTeamCUA and benchmark RTC-Bench, as well as the code running all evaluated CUAs.
@@
Appendix A RTC-Bench
A.1 Examples in RTC-Bench
To demonstrate the full scope of RTC-Bench, we provide detailed examples of indirect prompt injection tests in Sections  A.1.1,  A.1.2, and  A.1.3. These examples are showcased across our three available web platforms (RocketChat, Forum, OwnCloud) and represent our focused benign CUA use cases of System Configuration, Software Installation and Project Setup (described in § 2.1 and § 4.1). Our benign goals also are represented using two different instantiation types (General and Specific) to represent varying levels of user expertise, along with a Pointer instruction type, representing an Information-Gathering Assistant use case where CUAs just retrieve knowledge without execution (Appendix H.1), and a version for use in the End2End setting (see Appendix A.2 for more details). In addition, each example targets a distinct adversarial goal category within the CIA triad, representing fundamental security violations across the Confidentiality, Integrity, and Availability categories respectively. For high-resolution screenshots of indirect prompt injection (marked in red) within each web platform, see Figures 4, 5, 6. See later sections for additional examples and details about benign goals (A.2), adversarial scenarios (A.3), and adversarial examples (A.4).
A.1.1 Confidentiality
@@ A.3 Adversarial Scenarios
For our large-scale adversarial benchmark, we design a diverse set of adversarial scenarios that target the user’s local OS and define 24 total adversarial goals, each targeting a specific CIA security principle and corresponding to a distinct adversarial outcome (shown in Appendix A.4). For Confidentiality-based goals, we examine Web $\rightarrow$ OS $\rightarrow$ Web adversarial scenarios in which local OS files are exfiltrated to adversaries via external platforms, such as sending through a chat interface (RocketChat), uploading to shared storage (OwnCloud), or public posting content on online forums (Reddit). Meanwhile, adversarial goals for Integrity and Availability primarily target the data or functionality of the user’s local operating system for Web $\rightarrow$ OS attack pathways. For Availability, we focus on actions that degrade system performance, such as excessive CPU or disk usage, or disrupt services, such as disabling essential components like sshd. For file-based adversarial goals, we further classify target files into three categories to reflect varying harmful impacts to the user or system; Normal, representing generic, user-created file names (e.g. file.txt); Privacy Info, representing files directly related to sensitive user data (e.g. BankStatement.pdf); and System Info, representing critical system files directly related to core OS functionality (e.g. /etc/security). This categorization enables us to evaluate whether CUAs exhibit different levels of vulnerability depending on the sensitivity or impact of the targeted file.
A.4 Adversarial Examples
A.4.1 Confidentiality
@@ A.4.3 Availability
Appendix B RedTeamCUA Framework Diagram
Figure 7: Overview of our hybrid sandbox approach for systematic adversarial testing of CUAs. Built upon OSWorld (Xie et al., 2024; 2025), our sandbox integrates isolated web platforms to support realistic, end-to-end evaluation of adversarial scenarios spanning both web and OS interfaces simultaneously while preventing real-world harm. The Adversarial Task Initial State Config is used for flexible configuration of adversarial scenarios, defining adversarial injection content and locations, adversarial environment state initialization, and execution-based evaluators used to determine harmful task completion. For a clearer view of the three selected web platforms, high-resolution screenshots are provided in Appendix A.1.
Appendix C Comparison of CUA Evaluation Frameworks
Table 6: Comparison with previous evaluation frameworks that could be applied for adversarial testing of CUA across several key dimensions detailed in C. ‘–’ indicates cases that are not directly applicable or lack details in the original paper and $\sim$ represents cases where the framework has partial support for a specified dimension.
@@ Appendix C Comparison of CUA Evaluation Frameworks
Appendix D Experiment Setup Details
@@
Appendix E CIA Classification Principles
@@ Appendix F Licenses
Appendix G Defense Results
G.1 Prompt Injection Detection
@@
G.2 Defensive System Prompt
@@
We use the same 50 high-risk examples as in Appendix G.1 and evaluate Operator and Claude 3.7 Sonnet | CUA. As shown in Figure 8, the defensive system prompt serves as a simple and effective mitigation strategy and reduces ASR by nearly half compared to the default system prompt alone. However, the defensive system prompt alone is insufficient for secure deployment, as Claude 3.7 Sonnet | CUA and Operator (w/o checks) still achieve a high ASR around 50%. While this may be a valuable addition to the default configuration of CUAs, further research into more effective defensive system prompts (or defensive mechanisms in general) tailored to CUAs is still required for real-world deployment.
G.3 Model-level Defense
Beyond system-level defenses in G.1 and G.2, we also evaluate whether model-level defenses can mitigate risks in our setting. Meta SecAlign Chen et al. (2025d) is a recent open-source foundation model with built-in defenses against prompt injection, reportedly improving security in web navigation benchmarks such as WASP (Evtimov et al., 2025). We use Meta SecAlign 70B as the base model in our Adapted LLM-based Agent setting and evaluate it on the same 50 high-risk tasks. Since it is a text-only model, we use accessibility (a11y) tree as the agent observation modality.
@@ G.3 Model-level Defense
G.4 Discussions on Defenses for CUAs
@@
Appendix H Additional Results
H.1 Results by Benign and Adversarial Goal Type
@@
H.2 Results by Observation Type
@@
H.3 Results by File Type
@@
H.4 Utility (SR) Under Attack
@@ C I A C I A C I A
H.5 Attack Outcomes by Number of Attempts
@@ H.6 End2End Results Breakdown
Appendix I System Prompts

### v5

@@ 8 Conclusion
Acknowledgment
The authors would like to thank colleagues from the OSU NLP group for their constructive feedback. This research was sponsored in part by NSF CAREER #1942980, NSF CAREER #2443149, NSF CNS #2112471, the Alfred P. Sloan Foundation Fellowship, the Schmidt Sciences Safety Science award, and Ohio Supercomputer Center (Center, 1987). The views and conclusions contained herein are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notice herein.
@@ Acknowledgment
In § 3, we describe how we construct our hybrid sandbox RedTeamCUA, by leveraging OSWorld as a backbone and integrating Docker-based web replicas from WebArena and TheAgentCompany. In§ 4, we detail the formulation of benign and adversarial goals, along with concrete examples in Appendix B. In § 5.1 and Appendix E, we provide details on the evaluated CUAs, evaluation metrics (i.e., SR, ASR and AR) and AWS configuration. Upon acceptance, we will open-source all our related materials, including our sandbox RedTeamCUA and benchmark RTC-Bench, as well as the code running all evaluated CUAs.
@@
Appendix A Limitations
(3) Additionally, we observe that CUAs exhibit varying levels of vulnerability depending on the specific filename used (Appendix I.3). Although this observation alone does not justify specifying filenames in the injection, it highlights an additional consideration: filenames themselves might meaningfully influence attack outcomes, further reinforcing the need for controlled specification in our evaluations.
Appendix B RTC-Bench
B.1 Examples in RTC-Bench
To demonstrate the full scope of RTC-Bench, we provide detailed examples of indirect prompt injection tests in Sections  B.1.1,  B.1.2, and  B.1.3. These examples are showcased across our three available web platforms (RocketChat, Forum, OwnCloud) and represent our focused benign CUA use cases of System Configuration, Software Installation and Project Setup (described in § 2.1 and § 4.1). Our benign goals also are represented using two different instantiation types (General and Specific) to represent varying levels of user expertise, along with a Pointer instruction type, representing an Information-Gathering Assistant use case where CUAs just retrieve knowledge without execution (Appendix I.1), and a version for use in the End2End setting (see Appendix B.2 for more details). In addition, each example targets a distinct adversarial goal category within the CIA triad, representing fundamental security violations across the Confidentiality, Integrity, and Availability categories respectively. For high-resolution screenshots of indirect prompt injection (marked in red) within each web platform, see Figures 4, 5, 6. See later sections for additional examples and details about benign goals (B.2), adversarial scenarios (B.3), and adversarial examples (B.4).
B.1.1 Confidentiality
@@ B.3 Adversarial Scenarios
For our large-scale adversarial benchmark, we design a diverse set of adversarial scenarios that target the user’s local OS and define 24 total adversarial goals, each targeting a specific CIA security principle and corresponding to a distinct adversarial outcome (shown in Appendix B.4). For Confidentiality-based goals, we examine Web $\rightarrow$ OS $\rightarrow$ Web adversarial scenarios in which local OS files are exfiltrated to adversaries via external platforms, such as sending through a chat interface (RocketChat), uploading to shared storage (OwnCloud), or public posting content on online forums (Reddit). Meanwhile, adversarial goals for Integrity and Availability primarily target the data or functionality of the user’s local operating system for Web $\rightarrow$ OS attack pathways. For Availability, we focus on actions that degrade system performance, such as excessive CPU or disk usage, or disrupt services, such as disabling essential components like sshd. For file-based adversarial goals, we further classify target files into three categories to reflect varying harmful impacts to the user or system; Normal, representing generic, user-created file names (e.g. file.txt); Privacy Info, representing files directly related to sensitive user data (e.g. BankStatement.pdf); and System Info, representing critical system files directly related to core OS functionality (e.g. /etc/security). This categorization enables us to evaluate whether CUAs exhibit different levels of vulnerability depending on the sensitivity or impact of the targeted file.
B.4 Adversarial Examples
B.4.1 Confidentiality
@@ B.4.3 Availability
Appendix C RedTeamCUA Framework Diagram
Figure 7: Overview of our hybrid sandbox approach for systematic adversarial testing of CUAs. Built upon OSWorld (Xie et al., 2024; 2025), our sandbox integrates isolated web platforms to support realistic, end-to-end evaluation of adversarial scenarios spanning both web and OS interfaces simultaneously while preventing real-world harm. The Adversarial Task Initial State Config is used for flexible configuration of adversarial scenarios, defining adversarial injection content and locations, adversarial environment state initialization, and execution-based evaluators used to determine harmful task completion. For a clearer view of the three selected web platforms, high-resolution screenshots are provided in Appendix B.1.
Appendix D Comparison of CUA Evaluation Frameworks
Table 6: Comparison with previous evaluation frameworks that could be applied for adversarial testing of CUA across several key dimensions detailed in D. ‘–’ indicates cases that are not directly applicable or lack details in the original paper and $\sim$ represents cases where the framework has partial support for a specified dimension.
@@ Appendix D Comparison of CUA Evaluation Frameworks
Appendix E Experiment Setup Details
@@
Appendix F CIA Classification Principles
@@ Appendix G Licenses
Appendix H Defense Results
H.1 Prompt Injection Detection
@@
H.2 Defensive System Prompt
@@
We use the same 50 high-risk examples as in Appendix H.1 and evaluate Operator and Claude 3.7 Sonnet | CUA. As shown in Figure 8, the defensive system prompt serves as a simple and effective mitigation strategy and reduces ASR by nearly half compared to the default system prompt alone. However, the defensive system prompt alone is insufficient for secure deployment, as Claude 3.7 Sonnet | CUA and Operator (w/o checks) still achieve a high ASR around 50%. While this may be a valuable addition to the default configuration of CUAs, further research into more effective defensive system prompts (or defensive mechanisms in general) tailored to CUAs is still required for real-world deployment.
H.3 Model-level Defense
Beyond system-level defenses in H.1 and H.2, we also evaluate whether model-level defenses can mitigate risks in our setting. Meta SecAlign Chen et al. (2025d) is a recent open-source foundation model with built-in defenses against prompt injection, reportedly improving security in web navigation benchmarks such as WASP (Evtimov et al., 2025). We use Meta SecAlign 70B as the base model in our Adapted LLM-based Agent setting and evaluate it on the same 50 high-risk tasks. Since it is a text-only model, we use accessibility (a11y) tree as the agent observation modality.
@@ H.3 Model-level Defense
H.4 Discussions on Defenses for CUAs
@@
Appendix I Additional Results
I.1 Results by Benign and Adversarial Goal Type
@@
I.2 Results by Observation Type
@@
I.3 Results by File Type
@@
I.4 Utility (SR) Under Attack
@@ C I A C I A C I A
I.5 Attack Outcomes by Number of Attempts
@@ I.6 End2End Results Breakdown
Appendix J System Prompts